OpenStudy (anonymous):

The width of a door is 4 feet. The door and wall form a right angle. C= pi(d), pi= 22/7 What distance does the edge of the door move when the door is opened all the way? A. 4 feet B. 5 3/7 feet C. 6 2/7 feet D. 7 feet E. None of the above

OpenStudy (anonymous):

the door is 4

OpenStudy (anonymous):

whis the value of x x^2=4^2+4^2=4sqrt(2)

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

this what i think

OpenStudy (anonymous):

Thank you. :)